You invest $1,000.00 in one account and $3,000.00 in another account. Both accounts have the same interest rate over the same amount of time. How will the interest earned compare?
Answer by htmentor(1343)

In both cases, the amount of interest earned will be proportional to the principal.
In general, the interest is given by the formula
I = P((1+r)^n - 1), where P is the original principal, r is the interest rate, and t the term
Since r and n are the same, we can write
I = k*P, where k=constant
One account starts with $1000, so
I1 = 1000*k
The other account starts with $3000, so
I2 = 3000*k
The ratio of the interest earned is 3:1